OpenStudy (anonymous):

1) Marco and his two younger sisters would like to purchase a silver charm bracelet for their mother’s birthday. They went to the mall and found what they were looking for at Store A. In Store A, the bracelet without charms costs $85 and each charm costs $15. 2)If Marco and his sisters have saved $250, make a graph to show all of the charms they could purchase with a bracelet if they had saved between $85 and $250. (Use Geogebra to plot your function and zoom until you get the range between $85 and $250). Paste your function below:

OpenStudy (dan815):

ok

OpenStudy (dan815):

1)Cost of Bracelet = 85+15n, where n= number of charms they buy

OpenStudy (anonymous):

I need help with writing the equation to graph the 2nd one, can you help please :( ?

OpenStudy (dan815):

Okay im reading wait

OpenStudy (dan815):

The question is a bit confusing it says, they saved 250 then it says they saved 85 to 250... but either way, lets just graph the function

OpenStudy (dan815):

y=85+15n, lets graph for y=85 to y=250

OpenStudy (dan815):